Circular No. 7815

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2002 C2 (LINEAR) R. Huber and F. Shelley, Lincoln Laboratory, Massachusetts Institute

of Technology, report the discovery by LINEAR of a comet (discovery observation below) with a tail 42′′-47′′ long in p.a. 12^{o}.5. Cometary activity has been confirmed via CCD observations by A. Galád and L. Kornoš at Modra on Feb. 1.8 (m_{1} = 16.3, asymmetric coma extending 15′′ toward p.a. ∼ 30^{o}) and 2.7 UT (slightly diffuse with coma diameter ∼ 10′′), by J. Tichá and M. Tichý at Kleť on Feb. 2.7 (coma diameter 10′′), and by T. Payer at Düsseldorf, Germany, on Feb. 2.8 (coma visible, m_{1} = 16.6).

2002 UT  α_{2000}  δ_{2000}  m_{1}
Feb. 1.09074  2^{h}17^{m}07.^{s}09  +55^{o}10′02′′.3  17.8

Additional astrometry (Feb. 1–2), the following very preliminary parabolic orbital elements, and an ephemeris appear on MPEC 2002-C05.

T = 2002 Mar. 27.561 TT  ω = 156.^{o}393
Ω = 242.588 }2000.0
q = 3.28292 AU  i = 105.329

SUPERNOVA 2002ao IN UGC 9299 K. Kinugasa and H. Kawakita, Gunma Astronomical Observatory

(GAO); and H. Yamaoka, Kyushu University, report that a low-resolution spectrum of SN 2002ao (IAUC 7809, 7810) was obtained on Jan. 31.73 UT with the GAO 0.65-m telescope (+ GCS), showing a flat continuum with notable absorption features at 570 and 610 nm. The spectrum closely resembles that of type-Ic SN 1999bz near maximum light (IAUC 7162).

SUPERNOVA 2002H IN MCG –02-35-11 D. C. Leonard, University of Massachusetts at Amherst; and A. V.

Filippenko, University of California at Berkeley, report that inspection of a preliminarily reduced CCD spectrum (range 430–800 nm, resolution 0.7 nm), obtained with the Cerro Tololo 4-m telescope on Jan. 22.36 UT by L. E. Kay (Columbia University), H. D. Tran (Johns Hopkins University), and A. M. Magalhaes (Universidade de São Paulo), shows that SN 2002H (cf. IAUC 7798) is a type-Ia supernova ∼ 2–3 weeks past maximum light.

2002 February 2  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green